check whether every integrable function is differentiable or not ? justify
1
Expert's answer
2019-10-28T10:27:00-0400

Solution:


Every integrable function does not need to be differentiable.


A piecewise function (with discontinuity) is integrable but not differentiable.


Example(1): "y = |x|" is integrable in the domain [-1, 1], like


"\int y dy =\int_{-1}^1 y dy = \int_{-1}^ 0 (-x) dx + \int _0^1 x dx"

But this function |x| is not differentiable at x = 0.


Left hand Derivative



"lim_{x \to 0^-} \frac {f(x) - f{(0)} }{x - 0} =lim_{x \to 0^-} \frac {|x| - 0} {x - 0} = lim_ {x \to 0^-} \frac {-x} {x} \\
 = lim_{x \to 0^-} (-1) = - 1"



Right hand Derivative


"lim_{x \to 0^+} \frac {f(x) - f{(0)} }{x - 0} = lim_{x \to 0^+} \frac {|x| - 0} {x - 0} = lim_ {x \to 0^+} \frac {x} {x} \\
 = lim_{x \to 0^-} (1) = 1"

Left hand Derivative is not equal to Right hand  Derivative,

So, it is not differentiable.


Example(2): A step function is integrable but not differentiable


Example(3): A Weierstrass function is integrable but not differentiable.